OpenStudy (anonymous):

10+9+8+7+6+5+4+3+2+1

OpenStudy (anonymous):

55 is answer

OpenStudy (ankitshaw):

n(n+1)/2 will give u the answer where n is the last number

OpenStudy (ankitshaw):

u can use my eqn for sum of any no of natural nos

geerky42 (geerky42):

\[\sum_{i=1}^{n}i = n + (n-1) + (n-2) + ... + 3+2+1 = \dfrac{n(n+1)}{2}\]

geerky42 (geerky42):

Like what @ankitshaw said.

OpenStudy (ankitshaw):

@geerky42 thnax for that derivation its very imp to understand how a formula is derived

geerky42 (geerky42):

np @ankitshaw Is that clear? @ikatouni

OpenStudy (anonymous):

thanks guys